### v2

@@ 1 Introduction
Wireless capsule endoscopy (WCE) allows for effective observation of the internal structures of the gastrointestinal tract while minimizing patient discomfort ref1 . However, due to the limited illumination power of WCE and the complexity of the gastrointestinal environment, WCE images are often unclear, especially in the small intestine covered with tiny villi. An enlarged section of the real structure of the human small intestine is shown in Fig. 1 , revealing the surface of the small intestine covered with tiny villi responsible for nutrient absorption. Fig. 2 displays an actual image captured by WCE. A comparison shows that WCE images cannot effectively display the small intestinal villi structure and may even make it disappear, which is a common issue with WCE. The morphology of small intestinal villi serves as a crucial basis for diagnosing diseases such as indigestion and disorders in nutrient absorption. Enhancing techniques to highlight the microscopic villi structure and its edge details can assist doctors in better relying on villi morphology to observe small intestinal lesions. This enables a more precise identification of the location and severity of lesions, providing a more personalized treatment plan for patients ref2 . To the best of our knowledge, there are no articles in the literature reporting methods for enhancing WCE images to make villi more clearly visible.
@@
So far, there have been numerous related methods for enhancing small intestine images. The commonly used classic methods can be classified into three categories: Histogram Equalization (HE) methods ref4 ; ref5 ; ref6 ; ref16 , Retine-based methods ref7 ; ref8 ; ref9 ; ref17 ; ref18 and Unsharping mask (USM) methodsref10 ; ref11 ; ref12 ; ref13 ; ref14 ; ref19 ; ref20 , etc. Among them, the USM methods are the most advantageous approach in emphasizing image detail information. Methods based on HE enhance images by remapping the input image’s gray levels using the probability distribution of grayscale levels. Nonetheless, these methods encounter problems like inadequate enhancement, excessive enhancement, and notable noise amplification in the resulting enhancements. The Retinex-based methods view an image as a combination of light and reflection components, enhancing the image by adjusting the corresponding components. However, the effect of highlighting image detail information using the Retinex-based methods is not very ideal. Specifically, traditional USM methods use a fixed gain factor to enhance high-frequency components. Subsequently, some scholars improved the USM methods by proposing adaptive gain factors based on local image information ref21 . Clearly, the aforementioned methods were not specifically designed for enhancing small intestinal villi. Moreover, our experiments have revealed ref22 that they also do not achieve the desired effect of highlighting small intestinal villi. In addition, in the past decade, many scholars have proposed methods for endoscope image enhancement based on USM ref11 ; ref14 ; ref15 . While USM methods have the potential to emphasize image details, these methods have not adequately considered the balance between detail enhancement and noise suppression. This imbalance results in issues such as edge overshooting and amplified noise in the enhanced results. Therefore, they are not directly suitable for the purpose of enhancing small intestine villi in endoscopic images.
@@ 2 RELATED WORK
Three prevalent types of image enhancement methods include HE methods ref4 ; ref5 ; ref6 ; ref16 , Retine-based methodsref7 ; ref8 ; ref9 ; ref17 ; ref18 , and USM methodsref10 ; ref11 ; ref12 ; ref13 ; ref14 ; ref19 ; ref20 .
@@
The HE methodsref4 partition the image histogram by identifying local minima and assigning specific grayscale ranges to each partition. Subsequently, these partitions are equalized separately to enhance the image. Reference ref17 , the HE methods are improved by introducing brightness constraints to achieve image detail enhancement while preserving brightness. However, methods based on HE typically lead to insufficient or excessive enhancement of fine details, making them unsuitable for highlighting the villi in small intestine images.
@@
The Retinex-based methods view a scene image as a product of reflection and illumination components in the human eye ref18 , enhancing the scene image by adjusting these components accordinglyref3 . Reference ref19 , the decomposed reflection component is directly regarded as the enhancement result, but it can lead to the occurrence of halos and a loss of naturalness. Afterwards, scholars proposed a variational modelref10 for estimating reflection and illumination components, adjusting and combining these components to reconstruct the enhanced effect. While the Retinex-based methods’ effectiveness on fine details of microscopic villi structures is limited.
@@
The USM methods primarily enhance image details and edge information by emphasizing the image’s high-frequency content. Reference ref20 , the Laplacian operator is utilized for image filtering to enhance the image, but this method has many parameters and involves heavy computation. Reference ref21 , the difference between the iterative median filtering and the original image is used to obtain details of the image. However, this approach results in halo artifacts along the edges, with limited enhancement in smoothly lit areas. Reference ref11 , a Gaussian filter is applied to WCE images to accentuate image details, but this process can lead to issues like overshooting and excessive noise amplification. Referenceref12 , a hybrid median filterref23 calculates the median values for square, cross-shaped, and diagonal windows to replace the central pixel of the filtering window. However, this literature does not take into account the impact of noise during image enhancement. Reference ref13 , the high-frequency components of a guided image are added to the low-frequency components of the original image under the control of a gain factor. This factor is obtained using an optimization function based on guided filteringref24 and weighted guided filteringref25 . Additionally, in ref13 introduces a convolutional neural network to derive the gain factor. However, due to the considerable difficulty in obtaining in vivo data, this method is not applicable to images in an in-body environment. Although these methods improve upon USM techniques by enhancing filters and gain factors to boost detail enhancement, they still suffer from issues like edge overshooting and noise amplification when applied to image enhancement. The main problem lies in how gain factors are configured.
@@
To address this, this paper employs guided filteringref24 to filter WCE small intestine villi images, obtaining low-frequency components with well-preserved edges and gradients. Constructing light gain functions based on the low-frequency components of different regions in WCE small intestine villi images to adaptively generate light gain factors. Furthermore, we create adaptive gradient gain factors by building gradient gain functions using the Laplacian operator’s convolution results in various areas of WCE small intestine villi images. Ultimately, we combine the previously mentioned illumination gain factors and the gradient gain factors to create the adaptive gain coefficients needed for enhancing the Unsharp Masking (USM) method, thereby improving the USM approach. Our method effectively enhances the clarity of small intestine villi details while suppressing noise in darker regions and preventing edge overshooting.
@@ 3.1 Low-High Frequency Component Decomposition
As can be observed, the essence of the USM technique involves subtracting the low-frequency component from the original image to obtain the high-frequency component ref16 . Then, the high-frequency component is amplified by multiplying it with a gain factor. Finally, the amplified high-frequency component is superimposed on the original image to enhance the image’s edges. To achieve a more accurate high-frequency component, it is crucial to select a suitable low-pass filter that effectively preserves edges while capturing low-frequency components.
The guided filteringref24 exhibits a linear relationship between the guided image and the output image as follows.
@@ 4.2 Evaluation Metrics
Compared with natural images, WCE images have limited contrast, and the distribution of their intensities in the whole dynamic range is not uniform. Therefore, the IRMLE proposed in ref11 is used to evaluate the richness of detail information of WCE images, and the larger the value, the better the detail enhancement effect. The IRMLE expression is shown as follows.
@@
NIQEref26 is the no-reference evaluation image assessment metric, and the smaller its value, the better the quality. Its quality is expressed as the distance between the NSS feature model and the MVG fitted to the features extracted from the test image, and its expression is shown in refer to (11).
@@ 4.4 Experimental Comparison Analysis
In this paper, 50 WCE small intestine images were enhanced using HCUM ref11 , NLUMref12 , UMGFref13 and the method of this paper. Table 2 shows the mean values of PSNR, IRMLE and NIQE after processing by the above methods, where the best objective values are shown in bold. Fig. 11(a), (b), and(c) shows the PSNR, IRMLE and NIQE values for the WCE small intestine images after each method individually processes them. In Table 2, although HCUM shows a decent IRMLE value, its PSNR and NIQE values are poor, indicating that using this method to enhance the details of WCE small intestine images may result in excessive noise enhancement. Additionally, the structure of the enhanced WCE small intestine images is altered. In Table 2, the IRMLE value of UMGF decreased compared to the original image. The WCE small intestine villus images enhanced using this method exhibit excessive brightness enhancement, resulting in detail enhancement occurring in excessively bright and excessively dark areas. Fig. 12(b), (g), and (l) shows the results after HCUM enhancement, Fig. 12(d), (i), and (n) shows the results after UMGF enhancement. We can observe that the WCE small intestine villi images enhanced by HCUM and UMGF exhibit excessive noise enhancement, structural changes, and edge overshooting. In Table 2, NLUM exhibits good PSNR, IRMLE, and NIQE values. Additionally, the NLUM enhancement results in Fig. 12(c), (h), and (m) reveal that the structure of the WCE small intestine villi image is well-preserved without the occurrence of edge overshooting. However, there is an issue of excessive noise enhancement in the dark areas of the WCE small intestine image after enhancement.

### v3

@@ 1 Introduction
Wireless capsule endoscopy (WCE) allows for effective observation of the internal structures of the gastrointestinal tract while minimizing patient discomfort [1]. However, due to the limited illumination power of WCE and the complexity of the gastrointestinal environment, WCE images are often unclear, especially in the small intestine covered with tiny villi. An enlarged section of the real structure of the human small intestine is shown in Fig. 1 , revealing the surface of the small intestine covered with tiny villi responsible for nutrient absorption. Fig. 2 displays an actual image captured by WCE. A comparison shows that WCE images cannot effectively display the small intestinal villi structure and may even make it disappear, which is a common issue with WCE. The morphology of small intestinal villi serves as a crucial basis for diagnosing diseases such as indigestion and disorders in nutrient absorption. Enhancing techniques to highlight the microscopic villi structure and its edge details can assist doctors in better relying on villi morphology to observe small intestinal lesions. This enables a more precise identification of the location and severity of lesions, providing a more personalized treatment plan for patients [2]. To the best of our knowledge, there are no articles in the literature reporting methods for enhancing WCE images to make villi more clearly visible.
@@
So far, there have been numerous related methods for enhancing small intestine images. The commonly used classic methods can be classified into three categories: Histogram Equalization (HE) methods [4, 5, 6, 16], Retine-based methods [7, 8, 9, 17, 18] and Unsharping mask (USM) methods[10, 11, 12, 13, 14, 19, 20], etc. Among them, the USM methods are the most advantageous approach in emphasizing image detail information. Methods based on HE enhance images by remapping the input image’s gray levels using the probability distribution of grayscale levels. Nonetheless, these methods encounter problems like inadequate enhancement, excessive enhancement, and notable noise amplification in the resulting enhancements. The Retinex-based methods view an image as a combination of light and reflection components, enhancing the image by adjusting the corresponding components. However, the effect of highlighting image detail information using the Retinex-based methods is not very ideal. Specifically, traditional USM methods use a fixed gain factor to enhance high-frequency components. Subsequently, some scholars improved the USM methods by proposing adaptive gain factors based on local image information [21]. Clearly, the aforementioned methods were not specifically designed for enhancing small intestinal villi. Moreover, our experiments have revealed [22] that they also do not achieve the desired effect of highlighting small intestinal villi. In addition, in the past decade, many scholars have proposed methods for endoscope image enhancement based on USM [11, 14, 15]. While USM methods have the potential to emphasize image details, these methods have not adequately considered the balance between detail enhancement and noise suppression. This imbalance results in issues such as edge overshooting and amplified noise in the enhanced results. Therefore, they are not directly suitable for the purpose of enhancing small intestine villi in endoscopic images.
@@ 2 RELATED WORK
Three prevalent types of image enhancement methods include HE methods [4, 5, 6, 16], Retine-based methods[7, 8, 9, 17, 18], and USM methods[10, 11, 12, 13, 14, 19, 20].
@@
The HE methods[4] partition the image histogram by identifying local minima and assigning specific grayscale ranges to each partition. Subsequently, these partitions are equalized separately to enhance the image. Reference [17], the HE methods are improved by introducing brightness constraints to achieve image detail enhancement while preserving brightness. However, methods based on HE typically lead to insufficient or excessive enhancement of fine details, making them unsuitable for highlighting the villi in small intestine images.
@@
The Retinex-based methods view a scene image as a product of reflection and illumination components in the human eye [18], enhancing the scene image by adjusting these components accordingly[3]. Reference [19], the decomposed reflection component is directly regarded as the enhancement result, but it can lead to the occurrence of halos and a loss of naturalness. Afterwards, scholars proposed a variational model[10] for estimating reflection and illumination components, adjusting and combining these components to reconstruct the enhanced effect. While the Retinex-based methods’ effectiveness on fine details of microscopic villi structures is limited.
@@
The USM methods primarily enhance image details and edge information by emphasizing the image’s high-frequency content. Reference [20], the Laplacian operator is utilized for image filtering to enhance the image, but this method has many parameters and involves heavy computation. Reference [21], the difference between the iterative median filtering and the original image is used to obtain details of the image. However, this approach results in halo artifacts along the edges, with limited enhancement in smoothly lit areas. Reference [11], a Gaussian filter is applied to WCE images to accentuate image details, but this process can lead to issues like overshooting and excessive noise amplification. Reference[12], a hybrid median filter[23] calculates the median values for square, cross-shaped, and diagonal windows to replace the central pixel of the filtering window. However, this literature does not take into account the impact of noise during image enhancement. Reference [13], the high-frequency components of a guided image are added to the low-frequency components of the original image under the control of a gain factor. This factor is obtained using an optimization function based on guided filtering[24] and weighted guided filtering[25]. Additionally, in [13]introduces a convolutional neural network to derive the gain factor. However, due to the considerable difficulty in obtaining in vivo data, this method is not applicable to images in an in-body environment. Although these methods improve upon USM techniques by enhancing filters and gain factors to boost detail enhancement, they still suffer from issues like edge overshooting and noise amplification when applied to image enhancement. The main problem lies in how gain factors are configured.
@@
To address this, this paper employs guided filtering[24] to filter WCE small intestine villi images, obtaining low-frequency components with well-preserved edges and gradients. Constructing light gain functions based on the low-frequency components of different regions in WCE small intestine villi images to adaptively generate light gain factors. Furthermore, we create adaptive gradient gain factors by building gradient gain functions using the Laplacian operator’s convolution results in various areas of WCE small intestine villi images. Ultimately, we combine the previously mentioned illumination gain factors and the gradient gain factors to create the adaptive gain coefficients needed for enhancing the Unsharp Masking (USM) method, thereby improving the USM approach. Our method effectively enhances the clarity of small intestine villi details while suppressing noise in darker regions and preventing edge overshooting.
@@ 3.1 Low-High Frequency Component Decomposition
As can be observed, the essence of the USM technique involves subtracting the low-frequency component from the original image to obtain the high-frequency component [16]. Then, the high-frequency component is amplified by multiplying it with a gain factor. Finally, the amplified high-frequency component is superimposed on the original image to enhance the image’s edges. To achieve a more accurate high-frequency component, it is crucial to select a suitable low-pass filter that effectively preserves edges while capturing low-frequency components.
The guided filtering[24] exhibits a linear relationship between the guided image and the output image as follows.
@@ 4.2 Evaluation Metrics
Compared with natural images, WCE images have limited contrast, and the distribution of their intensities in the whole dynamic range is not uniform. Therefore, the IRMLE proposed in [11] is used to evaluate the richness of detail information of WCE images, and the larger the value, the better the detail enhancement effect. The IRMLE expression is shown as follows.
@@
NIQE[26]is the no-reference evaluation image assessment metric, and the smaller its value, the better the quality. Its quality is expressed as the distance between the NSS feature model and the MVG fitted to the features extracted from the test image, and its expression is shown in refer to (11).
@@ 4.4 Experimental Comparison Analysis
In this paper, 50 WCE small intestine images were enhanced using HCUM [11], NLUM[12], UMGF[13] and the method of this paper. Table 2 shows the mean values of PSNR, IRMLE and NIQE after processing by the above methods, where the best objective values are shown in bold. Fig. 11(a), (b), and(c) shows the PSNR, IRMLE and NIQE values for the WCE small intestine images after each method individually processes them. In Table 2, although HCUM shows a decent IRMLE value, its PSNR and NIQE values are poor, indicating that using this method to enhance the details of WCE small intestine images may result in excessive noise enhancement. Additionally, the structure of the enhanced WCE small intestine images is altered. In Table 2, the IRMLE value of UMGF decreased compared to the original image. The WCE small intestine villus images enhanced using this method exhibit excessive brightness enhancement, resulting in detail enhancement occurring in excessively bright and excessively dark areas. Fig. 12(b), (g), and (l) shows the results after HCUM enhancement, Fig. 12(d), (i), and (n) shows the results after UMGF enhancement. We can observe that the WCE small intestine villi images enhanced by HCUM and UMGF exhibit excessive noise enhancement, structural changes, and edge overshooting. In Table 2, NLUM exhibits good PSNR, IRMLE, and NIQE values. Additionally, the NLUM enhancement results in Fig. 12(c), (h), and (m) reveal that the structure of the WCE small intestine villi image is well-preserved without the occurrence of edge overshooting. However, there is an issue of excessive noise enhancement in the dark areas of the WCE small intestine image after enhancement.